# Conformally symplectic structures and the Lefschetz condition

### 18 Mar 2024 v1

@@ Abstract.
Symplectic manifolds certainly can have locally conformally symplectic structures which are not globally conformally symplectic, c.f Example 3 below (we also refer the reader to [3, 4, 1]). A sufficient condition for the desired implication, in the presence of a compatible almost complex structure, is the hard Lefschetz condition in degree $1$.
@@
A proof that the first condition implies the second is given in  [14, Proposition 3.3], which we streamline below, using the same essential idea. We note that in  [19, Theorem 5.4], Tomassini and Wang show that the second condition implies first (in fact, in all degrees). Namely, for an almost-Kähler manifold, the space ${\mathcal{H}}_{d}$ satisfies the hard Lefschetz condition whenever ${\mathcal{H}}_{d}={\mathcal{H}}_{d^{c}}$ (in all degrees). The remaining statements are included to give additional characterizations in terms of decompositions of $d$-harmonic forms.
@@
Finally, by  [8, Theorem 5.1], the last condition implies the first, since the space ¯${\mathcal{H}}_{\bar{\partial}}\cap{\mathcal{H}}_{\mu}$ always has the Lefschetz $\mathfrak{sl}(2)$-representation, and so satisfies hard Lefschetz on $H^{1}_{d}(M)\cong{\mathcal{H}}^{1}_{d}(M)$ if ¯${\mathcal{H}}^{1}_{\bar{\partial}}\cap{\mathcal{H}}^{1}_{\mu}={\mathcal{H}}^{1}_{d}(M)$. ∎
@@
A locally conformally symplectic structure $(\eta,\theta)$ on a symplectic manifold $(M,\omega)$ can indeed fail to be globally conformally symplectic if $\omega$ and $\eta$ do not share a compatible $J$. For example, the nilmanifold with structure equations $dx_{1}=x_{1}x_{3}$, $dx_{2}=-x_{2}x_{3}$, and $dx_{3}=dx_{4}=0$, with symplectic form $\omega=x_{1}x_{2}+x_{3}x_{4}$ has $(\theta=x_{3},\eta=x_{1}x_{4}+x_{2}x_{3})$ which is locally conformally symplectic, but not globally conformally symplectic. Note that $\omega$ does satisfy the hard Lefschetz in degree $1$.
On the other hand, the nilmanifold with structure equations $dx_{i}=x_{i-1}x_{1}$ for $i=3,4$, has locally conformally symplectic data ($\theta=x_{2},\eta=x_{1}x_{3}+x_{2}x_{4})$, which is not globally conformally symplectic, since $x_{2}\in H^{1}_{d}$ is non-trivial. Note that $\omega=x_{1}x_{4}+x_{2}x_{3}$ is symplectic, but does not satisfy hard Lefschetz in degree $1$. In this example, $\omega$ and $\eta$ do not share a compatible $J$.
We do not know of an example where $\omega$ is symplectic, $\eta$ is locally conformally symplectic but not globally conformally symplectic, while $\eta$ and $\omega$ share a compatible $J$. By Theorem 1, any such example could not satisfy the hard Lefschetz condition in degree $1$, and by Theorem 7 below, there are no examples on solvmanifolds.
@@
By hypothesis, $\omega$ satisfies hard Lefschetz in degree $1$, and therefore all degrees. By  [18, Theorem 4.2], any other symplectic form satisfies hard Lefschetz as well, so the corollary follows then from Theorem 1. ∎
@@
It follows from [5, 6] that any symplectic form on a $2$-dimensional complex torus or a hyperelliptic surface satisfies hard Lefschetz. We also refer the reader to [13, Proposition 4.10] where it is proved that the cohomology Kähler cone coincides with the symplectic cone on a $2$-dimensional complex torus (see also [11]). The corollary then follows from Theorem 1. ∎
@@
It follows from [18, Theorem 5.1, Theorem 6.1, Remark 6.2] that any symplectic form on $M$ satisfies hard Lefschetz. The corollary then follows from Theorem 1. ∎
The next theorem shows in particular that on compact solvmanifolds $\Gamma\backslash G$ with left-invariant structures, the hard Lefschetz assumption of Theorem 1 can be dropped. We recall that in such cases, the existence of a uniform lattice $\Gamma$ in a simply connected Lie group $G$ implies $G$ is unimodular [15], and if $\Gamma\backslash G$ is symplectic, then $G$ is solvable [7].
Theorem 7.
Let $\mathfrak{g}$ be a $2n$-dimensional unimodular Lie algebra with $\omega\in\bigwedge^{2}\mathfrak{g}^{*}$ satisfying $d\omega=0$. If $\eta\in\bigwedge^{2}\mathfrak{g}^{*}$, and $\omega$ and $\eta$ share a compatible $J$ on $\mathfrak{g}$, then $d\eta=0$.
In particular if $(\eta,\theta)$ is a (linear) locally conformally symplectic structure, and $\omega$ and $\eta$ share a compatible $J$, then $\eta$ is a symplectic.
Note the last example in Example 3 is unimodular, so the assumption that $\omega$ and $\eta$ share a compatible $J$ is essential.
Proof.
We denote by $h$ the almost-Kähler metric induced by $(\omega,J)$ and by $g$ the almost-Hermitian metric induced by $(\eta,J)$. Then, we can diagonalize $g$ with respect to the metric $h$ so there exists a $h$-orthonormal basis $\{e^{1},e^{2}=Je^{1},\ldots,e^{2n-1},e^{2n}=Je^{2n-1}\}$ of the dual of $\mathfrak{g}$ so that
$g=\sum_{i=1}^{n}\lambda_{i}\left(e^{2i-1}\otimes e^{2i-1}+e^{2i}\otimes e^{2i}\right),$
and so
$\eta=\sum_{i=1}^{n}\lambda_{i}\left(e^{2i-1}\wedge e^{2i}\right),$
where $\lambda_{i}$ are constant. On the other hand $\omega=\sum_{i=1}^{n}e^{2i-1}\wedge e^{2i}$. Hence,
$\sum_{i=1}^{n}d\left(e^{2i-1}\wedge e^{2i}\right)=0.$
The unimodular condition will be shown to imply that each summand above vanishes. Denote by $\{e_{i}\}$ the basis of $\mathfrak{g}$. We compute
$d\left(e^{2i-1}\wedge e^{2i}\right)(e_{2i-1},e_{2i},e_{2j-1}),\quad d\left(e^{2j-1}\wedge e^{2j}\right)(e_{2i-1},e_{2i},e_{2j-1}),$
$d\left(e^{2i-1}\wedge e^{2i}\right)(e_{2i-1},e_{2i},e_{2j}),\quad d\left(e^{2j-1}\wedge e^{2j}\right)(e_{2i-1},e_{2i},e_{2j}),$
for all indices $1\leq i,j\leq n$ with $i\neq j.$ For instance, for $i\neq j$,
(1) $d\left(e^{2i-1}\wedge e^{2i}\right)(e_{2i-1},e_{2i},e_{2j-1})=e^{2i-1}\left([e_{2i-1},e_{2j-1}]\right)+e^{2i}\left([e_{2i},e_{2j-1}]\right),$
and
(2) $d\left(e^{2j-1}\wedge e^{2j}\right)(e_{2i-1},e_{2i},e_{2j-1})=e^{2j}\left([e_{2i-1},e_{2i}]\right).$
Since $d\omega=0,$ we get from Equations (1) and (2) that
$e^{2i-1}\left([e_{2i-1},e_{2j-1}]\right)+e^{2i}\left([e_{2i},e_{2j-1}]\right)+e^{2j}\left([e_{2i-1},e_{2i}]\right)=0.$
Fixing $j$ and summing over $1\leq i\leq n$ with $i\neq j$, we get
$\sum_{i\neq j}e^{2i-1}\left([e_{2i-1},e_{2j-1}]\right)+\sum_{i\neq j}e^{2i}\left([e_{2i},e_{2j-1}]\right)+\sum_{i\neq j}e^{2j}\left([e_{2i-1},e_{2i}]\right)=0.$
Using the fact that $\mathfrak{g}$ is unimodular,
$e^{2j}\left([e_{2j-1},e_{2j}]\right)+\sum_{i\neq j}e^{2j}\left([e_{2i-1},e_{2i}]\right)=0,$
so that
$e^{2j}\left(\sum_{i=1}^{n}[e_{2i-1},e_{2i}]\right)=\sum_{i=1}^{n}e^{2j}\left([e_{2i-1},e_{2i}]\right)=0,$
for all $1\leq j\leq n$. Equivalently, for all $1\leq j\leq n$
$e^{2j-1}\left(\sum_{i=1}^{n}[e_{2i-1},e_{2i}]\right)=0.$
It follows that for each $i$, $\operatorname{span}\{e_{2i-1},e_{2i}\}$ is a Lie subalgebra of $\mathfrak{g}$ and therefore $d\left(e^{2i-1}\wedge e^{2i}\right)=0$ for all $i=1,\ldots,n$. Therefore, $\eta$ is a symplectic form. ∎
In the non-invariant case, the follow question alluded to in Example 3 seems to be open:
Question 8.
Is there a compact symplectic manifold $(M,\omega)$ with a locally conformally symplectic structure $(\eta,\theta)$, which is not globally conformally symplectic, such that $\eta$ and $\omega$ share a compatible $J$?
@@ Acknowledgements
The authors are grateful to Giovanni Bazzoni and Liviu Ornea for useful discussions about Question 8.

### 05 Apr 2024 v2

@@ Abstract.
Symplectic manifolds can have locally conformally symplectic structures which are not globally conformally symplectic, even in the presence of mutually compatible almost complex structures, c.f. Example 3 below (we also refer the reader to [3, 4, 1, 17]). A sufficient condition for the desired implication, in the presence of a compatible almost complex structure, is the hard Lefschetz condition in degree $1$.
@@
A proof that the first condition implies the second is given in  [13, Proposition 3.3], which we streamline below, using the same essential idea. We note that in  [18, Theorem 5.4], Tomassini and Wang show that the second condition implies first (in fact, in all degrees). Namely, for an almost-Kähler manifold, the space ${\mathcal{H}}_{d}$ satisfies the hard Lefschetz condition whenever ${\mathcal{H}}_{d}={\mathcal{H}}_{d^{c}}$ (in all degrees). The remaining statements are included to give additional characterizations in terms of decompositions of $d$-harmonic forms.
@@
Finally, by  [7, Theorem 5.1], the last condition implies the first, since the space ¯${\mathcal{H}}_{\bar{\partial}}\cap{\mathcal{H}}_{\mu}$ always has the Lefschetz $\mathfrak{sl}(2)$-representation, and so satisfies hard Lefschetz on $H^{1}_{d}(M)\cong{\mathcal{H}}^{1}_{d}(M)$ if ¯${\mathcal{H}}^{1}_{\bar{\partial}}\cap{\mathcal{H}}^{1}_{\mu}={\mathcal{H}}^{1}_{d}(M)$. ∎
@@
Two examples show the necessity of the assumptions in the theorem.
In [17, Example 4.1] the authors show that the Kodaira-Thurston manifold supports both a symplectic structure $\omega$ and a locally conformally symplectic structure $(\eta,\theta)$, which is not globally conformally symplectic, while $\eta$ and $\omega$ share a compatible $J$. In this case the hard Lefschetz condition fails in degree $1$.
On the other hand, a locally conformally symplectic structure $(\eta,\theta)$ on a symplectic manifold $(M,\omega)$ can fail to be globally conformally symplectic if $\omega$ and $\eta$ do not share a compatible $J$, even if the hard Lefschetz condition holds in degree $1$. For example, the nilmanifold with structure equations $dx_{1}=x_{1}x_{3}$, $dx_{2}=-x_{2}x_{3}$, and $dx_{3}=dx_{4}=0$, with symplectic form $\omega=x_{1}x_{2}+x_{3}x_{4}$ has $(\theta=-x_{3},\eta=x_{1}x_{4}+x_{2}x_{3})$ which is locally conformally symplectic, but not globally conformally symplectic.
@@
By hypothesis, $\omega$ satisfies hard Lefschetz in degree $1$, and therefore all degrees. By  [16, Theorem 4.2], any other symplectic form satisfies hard Lefschetz as well, so the corollary follows then from Theorem 1. ∎
@@
It follows from [5, 6] that any symplectic form on a $2$-dimensional complex torus or a hyperelliptic surface satisfies hard Lefschetz. We also refer the reader to [12, Proposition 4.10] where it is proved that the cohomology Kähler cone coincides with the symplectic cone on a $2$-dimensional complex torus (see also [10]). The corollary then follows from Theorem 1. ∎
@@
It follows from [16, Theorem 5.1, Theorem 6.1, Remark 6.2] that any symplectic form on $M$ satisfies hard Lefschetz. The corollary then follows from Theorem 1. ∎
@@ Acknowledgements
The authors are grateful to Adriano Tomassini and Nicolleta Tardini for pointing out their example in [17], and to Giovanni Bazzoni, Liviu Ornea, and Baptiste Chantraine for helpful discussions.